# Autonomous Shuttle Operation for Vulnerable Populations: Lessons and Experiences

### 27 Feb 2024 v1

@@ Abstract
The increasing shortage of drivers poses a significant threat to vulnerable populations, particularly seniors and disabled individuals who heavily depend on public transportation for accessing healthcare services and social events. Autonomous Vehicles (AVs) emerge as a promising alternative, offering potential improvements in accessibility and independence for these groups. However, current designs and studies often overlook the unique needs and experiences of these populations, leading to potential accessibility barriers. This paper presents a detailed case study of an autonomous shuttle test specifically tailored for seniors and disabled individuals, conducted during the early stages of the COVID-19 pandemic. The service, which lasted 13 weeks, catered to approximately 1,500 passengers in an urban setting, aiming to facilitate access to essential services. Drawing from the safety operator’s experiences and direct observations, we identify critical user experience and safety challenges faced by vulnerable passengers. Based on our findings, we propose targeted initiatives to enhance the safety, accessibility, and user education of AV technology for seniors and disabled individuals. These include increasing educational opportunities to familiarize these groups with AV technology, designing AVs with a focus on diversity and inclusion, and improving training programs for AV operators to address the unique needs of vulnerable populations. Through these initiatives, we aim to bridge the gap in AV accessibility and ensure that these technologies benefit all members of society.
@@ I Introduction
Existing research has explored public acceptance, trust, and behavior towards autonomous shuttles through road tests, with notable experiments conducted in environments like Mcity. These studies provide valuable insights into the operational and engagement aspects of deploying autonomous shuttles. However, the relevance of findings from settings like Mcity to vulnerable populations is limited, as these areas are not typically frequented by seniors or disabled individuals. This paper aims to bridge this gap by detailing a new deployment of autonomous shuttles in communities with a significant presence of vulnerable residents, focusing on providing transitional services among essential locations such as hospitals. In summary, this paper makes the following contributions:
  • We present comprehensive details of the first public road test for Autonomous Vehicles (AVs) specifically aimed at enhancing transportation accessibility for seniors and disabled individuals. This test focuses on facilitating their access to healthcare services and social events, aiming to fill the gap in current transportation solutions.
@@
The remainder of this paper is structured as follows: Section II outlines the detailed plan and methodology of the autonomous shuttle test. Section III discusses the observations and data collected by the safety operator during the testing period. Section V introduces four research initiatives aimed at enhancing autonomous vehicle technology, derived from insights gained during the test. Finally, Section VI provides a conclusion to the paper, summarizing the key findings and contributions.

### 28 Feb 2024 v2

@@ Abstract
The increasing shortage of drivers poses a significant threat to vulnerable populations, particularly seniors and disabled individuals who heavily depend on public transportation for accessing healthcare services and social events. Autonomous Vehicles (AVs) emerge as a promising alternative, offering potential improvements in accessibility and independence for these groups. However, current designs and studies often overlook the unique needs and experiences of these populations, leading to potential accessibility barriers. This paper presents a detailed case study of an autonomous shuttle pilot specifically tailored for seniors and disabled individuals, conducted during the early stages of the COVID-19 pandemic. The service, which lasted 13 weeks, catered to approximately 1500 passengers in an urban setting, aiming to facilitate access to essential services. Drawing from the safety operator’s experiences and direct observations, we identify critical user experience and safety challenges faced by vulnerable passengers. Based on our findings, we propose targeted initiatives to enhance the safety, accessibility, and user education of AV technology for seniors and disabled individuals. These include increasing educational opportunities to familiarize these groups with AV technology, designing AVs with a focus on diversity and inclusion, and improving training programs for AV operators to address the unique needs of vulnerable populations. Through these initiatives, we aim to bridge the gap in AV accessibility and ensure that these technologies benefit all members of society.
@@ I Introduction
Existing research has explored public acceptance, trust, and behavior towards autonomous shuttles through road pilots, with notable experiments conducted in environments like Mcity. These studies provide valuable insights into the operational and engagement aspects of deploying autonomous shuttles. However, the relevance of findings from settings like Mcity to vulnerable populations is limited, as these areas are not typically frequented by seniors or disabled individuals. This paper aims to bridge this gap by detailing a new deployment of autonomous shuttles in communities with a significant presence of vulnerable residents, focusing on providing transitional services among essential locations such as hospitals. In summary, this paper makes the following contributions:
  • We present comprehensive details of the first public road pilot for Autonomous Vehicles (AVs) specifically aimed at enhancing transportation accessibility for seniors and disabled individuals. This pilot focuses on facilitating their access to healthcare services and social events, aiming to fill the gap in current transportation solutions.
@@
The remainder of this paper is structured as follows: Section II outlines the detailed plan and methodology of the autonomous shuttle pilot. Section III discusses the observations and data collected by the safety operator during the testing period. Section V introduces four research initiatives aimed at enhancing autonomous vehicle technology, derived from insights gained during the pilot. Finally, Section VI provides a conclusion to the paper, summarizing the key findings and contributions.
@@ VI Summary
VII Acknowledgement
The authors wish to extend their sincere gratitude to Ray Smith from IXR Mobility for the opportunity to engage in the shuttle operation. This experience has been invaluable to our research and practical understanding in the autonomous driving field. We also express our appreciation to Navya Tech for providing essential training to our operators, which has been crucial for ensuring the safe implementation of our project activities.